Bramble Corporation is a small wholesaler of gourmet food products. Data regarding the store's operations follow:

  • Sales are budgeted at $340,000 for November, $320,000 for December, and $310,000 for January.
  • Collections are expected to be 80% in the month of sale and 20% in the month following the sale.
  • The cost of goods sold is 75% of sales.
  • The company would like to maintain ending merchandise inventories equal to 60% of the next month's cost of goods sold. Payment for merchandise is made in the month following the purchase.
  • Other monthly expenses to be paid in cash are $24,000.
  • Monthly depreciation is $15,000.
  • Ignore taxes.
Balance Sheet
October 31
Assets
Cash $ 20,000
Accounts receivable 70,000
Merchandise inventory 153,000
Property, plant and equipment, net of $572,000 accumulated depreciation 1,094,000
Total assets $ 1,337,000
Liabilities and Stockholders' Equity
Accounts payable $ 254,000
Common stock 820,000
Retained earnings 263,000
Total liabilities and stockholders' equity $ 1,337,000

The difference between cash receipts and cash disbursements for December would be:
Answer #1

Cash receipt for December = (340000*20%+320000*80%) = 324000

Purchase for November = (340000*75%)+(320000*75%*60%)-153000 = 246000

Cash payment for December = 246000+24000 = 270000

The difference between cash receipts and cash disbursement for December = 324000-270000 = 54000